OpenStudy (anonymous):

regarding indefinite integration, what would the antiderivetive of 5square root of x dx be?

OpenStudy (anonymous):

ok nevermind..

OpenStudy (anonymous):

\[\int5\sqrt{x}dx=5\int x^{1/2}dx=5\frac{x^{\frac{1}{2}+1}}{\frac{1}{2}+1}+C=5\frac{x^{3/2}}{3/2}+C\].

OpenStudy (roadjester):

You could probably simplify it one more step to \(\huge \frac {10} 3 x^{\frac 3 2 }+C\)

OpenStudy (nincompoop):

still not simplified

OpenStudy (wisp):

Completely simplified, it'd be: \[\frac{ 10 }{ 3 }\sqrt{x^{3}}+C\] Sure, you could pull the squared x out of the square root, but my teacher says she would accept this answer, so there's no need to waste the extra few seconds (like on the exam) to rewrite the whole step even further simplified. (P.S.) If you don't need additional help question anymore, you should state how to solve it (in case others look at it in the future for help) and close it ;).